OpenStudy (anonymous):

A computer hard disk starts from rest, then speeds up with an angular acceleration of 190rad/s2 until it reaches its final angular speed of 7200 rpm. How many revolutions has the disk made 10.0 s after it starts up?

OpenStudy (alekos):

have you attempted this question?

OpenStudy (anonymous):

Remember that the kinematics for rotational motion closely resemble the kinematics for linear motion just like displacement (s) and velocity (v) with linear kinematics \[s=v_it+\frac{1}{2}t^2\] \[s=vt\] and \[v=at\] for rotational motion, the angular displacement (θ) (number of radians traveled through) and angular velocity (ω) are given by \[ \theta = \omega_i t + \frac{1}{2}\alpha t^2\] \[\theta = \omega t\] and \[\omega=\alpha t\] Where alpha is the angular acceleration, and omega is the angular velocity. You need to find out how long it takes for the hard disk to reach a constant angular velocity, then add the angular displacement during acceleration to the angular displacement while it's angular velocity is constant (***hint: if it even reaches max speed in 10 seconds). You also need to remember that the angular displacement theta has to be converted into revolutions from radians \[1 \ \text{rev} = 2\pi \ \text{rad}\]